# Rec-RIR: Monaural Blind Room Impulse Response Identification via DNN-based Reverberant Speech Reconstruction in STFT Domain

### v1

@@ Abstract
Room impulse response (RIR) characterizes the complete propagation process of sound in an enclosed space. This paper presents Rec-RIR for monaural blind RIR identification. Rec-RIR is developed based on the convolutive transfer function (CTF) approximation, which models reverberation effect within narrow-band filter banks in the short-time Fourier transform (STFT) domain. Specifically, we propose a deep neural network (DNN) with cross-band and narrow-band blocks to estimate the CTF filter. The DNN is trained through reconstructing the noise-free reverberant speech spectra. This objective enables stable and straightforward supervised training. Subsequently, a pseudo intrusive measurement process is employed to convert the CTF filter estimate into time-domain RIR by simulating a common intrusive RIR measurement procedure. Experimental results demonstrate that Rec-RIR achieves state-of-the-art (SOTA) performance in both RIR identification and acoustic parameter estimation. Open-source codes are available online at https://github.com/Audio-WestlakeU/Rec-RIR.
@@
Index Terms—  System identification, room impulse response, acoustic parameters, convolutive transfer function, deep learning
1 Introduction
@@
Room impulse response (RIR) models the complete propagation process of sound in an enclosure and contains key information of acoustic characteristics. Accurate RIR estimation facilitates a wide range of applications, such as speech enhancement [1], speech recognition [2] and augmented/virtual reality [3]. A common intrusive method for measuring RIR is implemented by playing a known excitation signal and processing the measured recording signal [4]. However, intrusive RIR measurement is expensive and not always practical in real-world applications. In contrast, blind RIR identification has a broader range of applications. It aims to estimate RIRs solely from observed signals (e.g., reverberant speech signals) without known source signals (e.g., clean speech signals) [5, 6, 7].
@@
With the development of deep neural networks (DNNs), the performance of blind RIR identification has improved significantly. Many approaches encode fixed-length reverberant speech in the time domain and use a decoder to obtain fixed-length RIR estimates. S2IR-GAN [2] uses a generative adversarial network (GAN)-based architecture to estimate the RIR up to 0.25 s. The discriminator network is optimized to differentiate the estimated RIR from the ground truth RIR. Estimating RIR in the time domain remains challenging due to its long taps, which typically consist of thousands of samples. In FiNS [8], the RIR is decomposed into three distinct components: the direct-path impulse, early reflections and decaying filtered noise signals. A time-domain encoder-decoder architecture is employed to learn both the early RIR and the noise filter, which significantly reduces the output dimension, thus alleviating the difficulty of estimating long RIRs. SG-RIR [9] employs a segmental network that generates one segment of the complete RIR each time, with the network architecture shared across all segments. According to the convolutive transfer function (CTF) approximation [10], the reverberation effect can be modeled within narrow-band filter banks in the short-time Fourier transform (STFT) domain. Therefore, the estimation of long RIR can be replaced by estimating a much shorter CTF filter in the STFT domain, thus simplifying the task. In the STFT domain, BUDDy [11] and VINP [12] estimate RIRs by implicitly reconstructing the reverberant spectra. Neither of them directly estimates the filter through DNN. Instead, since BUDDy and VINP have additional tasks besides RIR identification, both of them iteratively optimize the estimation of clean speech and the filter in STFT domain based on the signal model and DNN predictions.
@@
This work presents Rec-RIR for monaural blind RIR identification. The basis of Rec-RIR is the CTF approximation, which models the reverberation effect within a CTF filter in the STFT domain. Rec-RIR comprises two main steps: CTF filter estimation and a pseudo intrusive measurement process. Using SpatialNet [13] and its derivative methods [12, 14, 15] as the backbone, we propose an end-to-end DNN consisting of cross-band and narrow-band blocks for CTF estimation. The proposed DNN outputs a fixed-length CTF filter based on the input of arbitrary length, allowing Rec-RIR to make full use of the long input recording. Since RIR is independent of noise, and reverberant speech is the filtered version of clean speech through RIR, we propose to learn the embeddings of reverberant speech and clean speech in the latent space, and use both to estimate the CTF filter. We use the reconstruction error of the reverberant spectrum as the primary objective for supervised training, and incorporate two auxiliary losses to provide guidance for the learning of embeddings. The objective enables stable and straightforward estimation of the CTF filter and, compared with BUDDy and VINP, allows the network to directly estimate the CTF filter without iteration. Further, a pseudo intrusive measurement process is introduced to convert the CTF filter into RIR by simulating a common intrusive RIR measurement procedure, as proposed in [12]. Experimental results demonstrate that Rec-RIR can accurately estimate RIRs. Additionally, it achieves state-of-the-art (SOTA) performance in the estimation of acoustic parameters, including reverberation time (RT60), direct-to-reverberant ratio (DRR), and clarity (C50).
The remainder of this paper is organized as follows: Section 2 formulates the signal model. Section 3 details the proposed Rec-RIR. Experiments and discussions are presented in Section 4. Finally, Section 5 concludes the entire paper.
2 Signal Model
@@
where $n$ is the sample index, $x(n)$ and $w(n)$ denote reverberant speech and additive noise, respectively. Reverberant speech $x(n)$ can be expressed as the convolution of clean speech $s(n)$ and RIR $h(n)$. For simplicity, we assume that the RIR starts exactly at the sample of the direct-path impulse.
Applying STFT to the time-domain signal model, according to the CTF approximation [10], we have
@@
3 Proposed Method
Rec-RIR aims to estimate RIRs given the noisy and reverberant speech recordings. Specifically, we propose a DNN to learn the mapping from the input spectrum $\mathbf{Y}$ to the CTF filter estimate $\hat{\mathbf{H}}$. Subsequently, a pseudo intrusive measurement process is introduced to convert the CTF filter estimate $\hat{\mathbf{H}}$ into RIR estimate $\hat{h}(n)$. The workflow of Rec-RIR is illustrated in Fig. 1.
Fig. 1: Workflow of Rec-RIR.
3.1 Proposed Network
3.1.1 Feature representation
The input and output features of the proposed DNN are constructed by concatenating the real and imaginary parts of the input spectrum $\mathbf{Y}$ and CTF filter $\hat{\mathbf{H}}$ along the expanded channel dimension. The input feature is
@@ 3.1.1 Feature representation
Fig. 2: Architecture of the proposed network.
3.1.2 Modules
@@
As illustrated in Fig. 2, the proposed DNN consists of an input module, a denoising module, a dereverberation module, and a CTF estimation module. The input and output sizes are labeled above and below each block, respectively. For brevity, the reshape and dimension permutation operations are omitted.
@@
Input module This module is designed for extracting the embeddings from the noisy and reverberant recording. It consists of a single 1-D convolutional layer applied along the frame axis, with a kernel size of 5 and $C$ channels, named as T-Conv1d. Through this input module, the 2-channel input features are expanded into $C$-dimensional embeddings for each time-frequency (T-F) bin.
@@
Denoising module and dereverberation module Since RIR is independent of noise, and reverberant speech is the filtered version of clean speech through RIR, we propose denoising and dereverberation modules to learn the embeddings of reverberant speech and clean speech, respectively. Both modules share the same network structure, which is derived from SpatialNet [13] and oSpatialNet [14] and has been proven effective for denoising and dereverberation tasks. Specifically, the cross-band block, composed of convolutional layers along the frequency axis and linear layers, processes all frames independently using identical model parameters, and remains consistent with the original design of SpatialNet. In general, RIR identification is an offline task. Therefore, we adopt the bidirectional Mamba-based narrow-band block proposed in [12], which consists of stacked one forward Mamba layer and one backward Mamba layer. The narrow-band block processes all frequency bands independently with the same model parameters. The denoising module and the dereverberation module contain $M_{\mathrm{noi}}$ and $M_{\mathrm{rev}}$ stacked cross-band and narrow-band blocks, respectively. Additionally, we employ two decoders that consist of two linear layers and LeakyReLU activation to transform these embeddings into reverberant and clean spectra. Decoders require the output embeddings to contain key information of reverberant and clean speech, providing direct guidance for training.
@@
CTF estimation module The outputs of the denoising and dereverberation modules are embeddings containing key information of reverberant and clean speech. The CTF estimation module takes both, and outputs the fixed-length CTF filter estimate. Specifically, the outputs of the denoising and dereverberation modules are weighted and summed to form the input of CTF estimation module with learnable weighting parameters $\alpha$ and $\beta$. According to the CTF approximation, the narrow-band filter banks across different frequency bands are independent. Therefore, $M_{\mathrm{ctf}}$ stacked narrow-band blocks are employed to estimate CTF embeddings. After that, the weight block, which consists of linear layers, LeakyReLU activation and softmax activation, is designed to calculate the importance of CTF embeddings at all frames. These weights are distinct for each frequency band. Within each band, the sum of weights along the frame axis equals 1. The weighting mechanism ensures a fixed-length output regardless of the input length, allowing Rec-RIR to accept inputs of arbitrary length and make full use of long input recordings. After weighting, we employ a CTF decoder to transform the CTF embeddings into the CTF filter. The CTF decoder shares the same structure as the aforementioned decoders, but differs in that it outputs $2L$ dimensions for each T-F bin.
3.1.3 Loss function
@@
We design a composite loss function for training the proposed network. It comprises three components: the reconstruction error of reverberant spectrum $\mathcal{L}_{\mathrm{rec}}$, the prediction error of reverberant spectrum $\mathcal{L}_{\mathrm{rvb}}$, and the prediction error of clean spectrum $\mathcal{L}_{\mathrm{cln}}$, which is
@@
where $\mathcal{L}_{\mathrm{rec}}=\mathcal{L}_{\mathrm{RI+Mag}}(\hat{\mathbf{H}}\circledast\mathbf{S},\mathbf{X})$, $\mathcal{L}_{\mathrm{rvb}}=\mathcal{L}_{\mathrm{RI+Mag}}(\hat{\mathbf{X}},\mathbf{X})$, $\mathcal{L}_{\mathrm{cln}}=\mathcal{L}_{\mathrm{RI+Mag}}(\hat{\mathbf{S}},\mathbf{S})$. $\lambda_{\mathrm{rvb}}$ and $\lambda_{\mathrm{cln}}$ are hyperparameters. $\mathcal{L}_{\mathrm{RI+Mag}}(\cdot,\cdot)$ is defined as [18]
$\mathcal{L}_{\mathrm{RI+Mag}}(\mathbf{X},\mathbf{Y})$ $=\frac{1}{FT}\left[||\mathbf{X}-\mathbf{Y}||_{1}+||\mathrm{Re}\{\mathbf{X}\}-\mathrm{Re}\{\mathbf{Y}\}||_{1}\right$. (7)
@@
Among these losses, $\mathcal{L}_{\mathrm{rec}}$ serves as the primary loss, guiding the proposed network to stably learn the direct mapping from the input spectrum to the CTF filter. $\mathcal{L}_{\mathrm{rvb}}$ and $\mathcal{L}_{\mathrm{cln}}$ act as auxiliary losses, helping the network obtain better embedding representations.
3.2 Pseudo intrusive measurement process
We employ a pseudo intrusive measurement process [12] to convert the CTF filter to RIR by simulating a common intrusive measurement process. Given a logarithmic sine sweep as an excitation signal $e(n)$, there is an inverse filter $v(n)$ that satisfies $e(n)*v(n)=\delta(n)$, where $\delta(n)$ is an impulse [4, 19]. Playing such an excitation signal, according to the signal model, the STFT of noise-free measurement signal is
$\mathbf{Z}\approx\hat{\mathbf{H}}\circledast\mathbf{E},$ (8)
where $\mathbf{E}$ is the STFT of excitation signal. The RIR is estimated by inverse filtering
$\hat{h}(n)=z(n)*v(n),$ (9)
where $z(n)$ is the inverse STFT of $\mathbf{Z}$.
4 Experiments
Table 1: Blind RIR identification results on SimACE.
@@ 4 Experiments
  FiNS [8] 0.097 0.113 0.167 0.857 2.153 2.639 0.737 6.489 6.597 0.925
BUDDy (pre-trained) [11] 0.054 0.124 0.173 0.941 3.931 4.570 0.783 4.290 4.654 0.773
BUDDy [11] 0.057 0.122 0.166 0.952 3.673 4.360 0.774 4.109 4.477 0.779
VINP-TCN+SA+S [12] 0.050 0.089 0.124 0.934 3.256 3.764 0.872 0.914 1.135 0.961
VINP-oSpatialNet [12] 0.050 0.103 0.154 0.895 2.398 2.875 0.908 0.977 1.295 0.943
  Rec-RIR 0.040 0.069 0.104 0.994 0.684 0.794 0.994 0.858 1.019 0.978
4.1 Datasets
We focus on audios sampled at 16 kHz. We use the same training and test sets as in VINP [12], which are generated by convolving clean speech with RIRs, followed by noise addition.

### v2

@@ Abstract
This paper presents Rec-RIR for monaural blind room impulse response (RIR) identification. Rec-RIR is developed based on the convolutive transfer function (CTF) approximation, which models reverberation effect within narrow-band filter banks in the short-time Fourier transform domain. Specifically, we propose a deep neural network (DNN) with cross-band and narrow-band blocks to estimate the CTF filter. The DNN is trained through reconstructing the noise-free reverberant speech spectra. This objective enables stable and straightforward supervised training. Subsequently, a pseudo intrusive measurement process is employed to convert the CTF filter estimate into RIR by simulating a common intrusive RIR measurement procedure. Experimental results demonstrate that Rec-RIR achieves state-of-the-art performance in both RIR identification and acoustic parameter estimation. Open-source codes are available online at https://github.com/Audio-WestlakeU/Rec-RIR.
Index Terms:
@@
System identification, room impulse response, acoustic parameters, convolutive transfer function, deep learning
I Introduction
@@
Room impulse response (RIR) models the complete propagation process of sound and contains key information of the acoustic environment. Accurate RIR estimation facilitates a wide range of applications, such as speech enhancement [28], speech recognition [17] and augmented/virtual reality [20]. A common intrusive method for measuring RIR is to play a known excitation signal, such as a maximum length sequence or a sine sweep signal, and then perform inverse filtering on the observation [22, 7]. However, intrusive measurement is expensive and not always practical in real-world applications. In contrast, blind RIR identification aims to estimate normalized RIRs solely from observation of a specific form (e.g., reverberant speech) without known source signals [11, 2, 3].
@@
Deep neural network (DNN)-based blind RIR identification use data-driven methods to learn normalized RIRs directly. Many DNN-based approaches encode fixed-length reverberant speech in the time domain and use a decoder to obtain fixed-length RIR estimates. S2IR-GAN [17] uses a generative adversarial network (GAN)-based architecture to estimate the RIR up to 0.25 s. The discriminator network is optimized to differentiate the estimated RIR from the ground truth RIR. Estimating RIR in the time domain remains challenging due to its long taps, which typically consist of thousands of samples. In FiNS [23], the RIR is decomposed into three distinct components: the direct-path impulse, early reflections and decaying filtered noise signals. A time-domain encoder-decoder architecture is employed to learn both the early RIR and the noise filter, which significantly reduces the output dimension, thus alleviating the difficulty of estimating long RIRs. SG-RIR [10] employs a segmental network that generates one segment of the complete RIR each time, with the network architecture shared across all segments. According to the convolutive transfer function (CTF) approximation [24], the reverberation effect can be modeled within narrow-band filter banks in the short-time Fourier transform (STFT) domain. Therefore, the estimation of long RIR can be replaced by estimating a much shorter CTF filter in the STFT domain, thus simplifying the task. BUDDy [9] and VINP [29] estimate RIRs by implicitly reconstructing the reverberant spectra. Both of them estimate the clean speech and the filter simultaneously and iteratively in STFT domain based on the signal model and DNN predictions.
@@
This work presents Rec-RIR for monaural blind RIR identification. The basis of Rec-RIR is the CTF approximation, which models the reverberation effect within narrow-band filtering. Rec-RIR comprises two main steps: CTF filter estimation and a pseudo intrusive measurement process. Using SpatialNet [16] and its derivative methods [29, 15, 21] as the backbone, we propose an end-to-end DNN consisting of cross-band and narrow-band blocks for CTF estimation. The proposed DNN outputs a fixed-length CTF filter based on the input of arbitrary length, allowing Rec-RIR to make full use of the long input recording. Since RIR is independent of noise, and reverberant speech is the filtered version of clean speech through RIR, we propose to learn the embeddings of reverberant speech and clean speech in the latent space, and use both to estimate the CTF filter. We use the reconstruction error of the reverberant spectrum as the primary objective for supervised training, and incorporate two auxiliary losses to provide guidance for the learning of embeddings. The objective enables stable and straightforward estimation of the CTF filter and, compared with BUDDy and VINP, allows the network to directly estimate the CTF filter without iteration. Further, a pseudo intrusive measurement process is introduced to convert the CTF filter into RIR by simulating a common intrusive RIR measurement procedure, as proposed in [29]. Experimental results demonstrate that Rec-RIR can accurately estimate RIRs and achieves state-of-the-art (SOTA) performance in the estimation of acoustic parameters, including reverberation time (RT60), direct-to-reverberant ratio (DRR), and clarity (C50).
II Signal Model
@@
where $n$ is the sample index, $x(n)$ and $w(n)$ denote reverberant speech and additive noise, respectively. Reverberant speech $x(n)$ can be expressed as the convolution of clean speech $s(n)$ and RIR $h(n)$. Notice that we constrain $s(n)$ to have the delay and gain of the direct-path observation, and constrain $h(n)$ to start with the direct-path impulse, to obtain a unique solution for RIR identification.
According to the CTF approximation [24], in the STFT domain, we have
@@
III Proposed Method
In Rec-RIR, we propose a DNN to learn the mapping from the input spectrum $\mathbf{Y}$ to the CTF filter estimate $\hat{\mathbf{H}}$. Subsequently, a pseudo intrusive measurement process is introduced to convert the CTF filter estimate $\hat{\mathbf{H}}$ into RIR estimate $\hat{h}(n)$. The workflow of Rec-RIR is illustrated in Fig. 1.
Figure 1: Workflow of Rec-RIR.
III-A Proposed Network
III-A1 Feature representation
The input and output features of the proposed DNN are constructed by concatenating the real and imaginary parts of the input spectrum $\mathbf{Y}$ and CTF filter $\hat{\mathbf{H}}$ along an expanded channel dimension. The input feature is
@@ III-A1 Feature representation
Figure 2: Architecture of the proposed network.
III-A2 Modules
@@
As shown in Fig. 2, the proposed DNN consists of an input module, a denoising module, a dereverberation module, and a CTF estimation module. The input and output sizes are labeled above and below each block, respectively. The reshape and dimension permutation operations are omitted.
@@
Input module This module extracts embeddings from noisy and reverberant recordings. It applies a single 1-D convolutional layer (T-Conv1d) with kernel size 5 and C channels along the frame axis, expanding 2-channel inputs into C-dimensional embeddings per time-frequency (T-F) bin, followed by PReLU activation.
@@
Denoising module and dereverberation module Since RIR is related to reverberation and clean speech but not to noise, we propose denoising and dereverberation modules to learn the embeddings of reverberant speech and clean speech, respectively. Both modules share the same network structure, which reuses the basic architecture from SpatialNet[16] and its extensions[29, 15], and has been proven effective for both denoising and dereverberation tasks. The cross-band block, composed of convolutional layers along the frequency axis and linear layers, processes all frames independently using identical model parameters, and remains consistent with the original design of SpatialNet. Note that as in SpatialNet, the ’F-Linear networks’ in cross-band blocks are shared within the denoising module and the dereverberation module, respectively. However, they are kept separate between these two modules. In general, RIR identification is an offline task. Therefore, we adopt the bidirectional Mamba-based narrow-band block proposed in [29], which consists of stacked one forward Mamba layer and one backward Mamba layer. The narrow-band block processes all frequency bands independently with the same model parameters. The denoising module and the dereverberation module contain $M_{\mathrm{noi}}$ and $M_{\mathrm{rev}}$ stacked cross-band and narrow-band blocks, respectively. To provide direct guidance for training, we employ two decoders that consist of two linear layers and LeakyReLU activation to transform the output embeddings of both module into reverberant and clean spectra respectively, which require the output embeddings to contain key information of reverberant and clean speech.
@@
CTF estimation module This module takes both reverberant and clean speech embeddings, and outputs the fixed-length CTF filter estimate. Specifically, the outputs of the denoising and dereverberation modules are weighted and summed to form the input of CTF estimation module with learnable weighting parameters $\alpha$ and $\beta$. According to the CTF approximation, the narrow-band filter banks across different frequency bands are independent. Therefore, $M_{\mathrm{ctf}}$ stacked narrow-band blocks are employed to estimate CTF embeddings. After that, a weight block, which consists of linear layers, LeakyReLU and Softmax activation, is designed to calculate the importance of CTF embeddings at all frames. These weights are distinct for each frequency band. Within each band, the sum of weights along the frame axis equals 1. The weighting mechanism ensures a fixed-length output regardless of the input length, allowing Rec-RIR to accept inputs of arbitrary length and make full use of long input recordings. After weighting, we employ a CTF decoder to transform the CTF embeddings into the CTF filter coefficients. The CTF decoder shares the same structure as the aforementioned decoders, but differs in that it outputs $2L$ dimensions for each T-F bin.
III-A3 Loss function
@@
We design a composite loss function for training, which comprises three components: the reconstruction error of reverberant spectrum $\mathcal{L}_{\mathrm{rec}}$, the prediction error of reverberant spectrum $\mathcal{L}_{\mathrm{rvb}}$, and the prediction error of clean spectrum $\mathcal{L}_{\mathrm{cln}}$, which is
@@
where $\mathcal{L}_{\mathrm{rec}}=\mathcal{L}_{\mathrm{RI+Mag}}(\hat{\mathbf{H}}\circledast\mathbf{S},\mathbf{X})$, $\mathcal{L}_{\mathrm{rvb}}=\mathcal{L}_{\mathrm{RI+Mag}}(\hat{\mathbf{X}},\mathbf{X})$, $\mathcal{L}_{\mathrm{cln}}=\mathcal{L}_{\mathrm{RI+Mag}}(\hat{\mathbf{S}},\mathbf{S})$. $\lambda_{\mathrm{rvb}}$ and $\lambda_{\mathrm{cln}}$ are hyperparameters. $\hat{\mathbf{H}}$, $\hat{\mathbf{X}}$ and $\hat{\mathbf{S}}$ are the output of decoders, shown in Fig. 2. $\mathcal{L}_{\mathrm{RI+Mag}}(\cdot,\cdot)$ is defined as [30]
$\mathcal{L}_{\mathrm{RI+Mag}}(\mathbf{X},\mathbf{Y})$ $=\frac{1}{FT}\left[|||\mathbf{X}|-|\mathbf{Y}|||_{1}+||\mathrm{Re}\{\mathbf{X}\}-\mathrm{Re}\{\mathbf{Y}\}||_{1}\right$. (7)
@@
Here, $\mathcal{L}_{\mathrm{rec}}$ serves as the primary loss, guiding the proposed network to stably learn the direct mapping from the input spectrum to the CTF filter. $\mathcal{L}_{\mathrm{rvb}}$ and $\mathcal{L}_{\mathrm{cln}}$ act as auxiliary losses, helping the network obtain better hidden representations.
III-B Pseudo intrusive measurement process
We employ the pseudo intrusive measurement process proposed in [29] to convert the CTF filter to RIR by simulating a common intrusive measurement process. Given a logarithmic sine sweep as an excitation signal $e(n)$, there is an inverse filter $v(n)$ that satisfies $e(n)*v(n)=\delta(n)$, where $\delta(n)$ is an impulse [22, 6]. Playing such an excitation signal, according to the signal model, the STFT of noise-free measurement signal is $\mathbf{Z}\approx\hat{\mathbf{H}}\circledast\mathbf{E}$, where $\mathbf{E}$ is the STFT of excitation signal. After all, the RIR is estimated by inverse filtering $\hat{h}(n)=z(n)*v(n)$, where $z(n)$ is the inverse STFT of $\mathbf{Z}$.
IV Experiments
TABLE I: Blind RIR identification results on SimACE.
@@ IV Experiments
  FiNS [23] 0.067 0.409 0.113 0.167 0.857 2.153 2.639 0.737 6.489 6.597 0.925
BUDDy (pre-trained) [9] 0.054 0.655 0.124 0.173 0.941 3.931 4.570 0.783 4.290 4.654 0.773
BUDDy [9] 0.057 0.621 0.122 0.166 0.952 3.673 4.360 0.774 4.109 4.477 0.779
VINP-TCN+SA+S [29] 0.050 0.695 0.089 0.124 0.934 3.256 3.764 0.872 0.914 1.135 0.961
VINP-oSpatialNet [29] 0.050 0.703 0.103 0.154 0.895 2.398 2.875 0.908 0.977 1.295 0.943
  Rec-RIR 0.040 0.805 0.069 0.104 0.994 0.684 0.794 0.994 0.858 1.019 0.978
IV-A Datasets
We use the same 16 kHz training and test sets as in VINP [29], which are generated by convolving clean speech with RIRs, followed by noise addition.